calculate the solubility of oxygen in water at 1.50 atm and 5°C if the solubility is 0.032 g oxygen per liter of water at 1.20 atm and 5°C


1
Expert's answer
2020-12-30T07:30:47-0500

Solubility = mass/molar mass × 1000/vol

Volume is known


But PV = nRT

n= 0.032g/32gmol-1 = 0.001mol


V = nRT/ P = 0.001mol × 0.08206atmkmol-1 × 278k / 1.20atm

V=0.019L


Solubility = mole × 1000/vol

= 0.001mol × 1000/0.019L

= 52.63mol/L